# Sparse Narrow-Band Topology Optimization for Large-Scale Thermal-Fluid Applications

### v1

@@ 4.1 Numerical method and preconditioning
Because the matrix $\mathbf{S}$ is symmetric positive definite, we solve (30) with the conjugate-gradient (CG) method, which can be viewed as a Krylov acceleration of the classical Uzawa algorithm. Each application of $\mathbf{S}$ requires the action of the inverse of the perturbed velocity Laplacian $\nu\overline{\mathbf{L}}_{U}+\overline{\mathbf{A}}(\Gamma)$. Forming this inverse explicitly would produce a dense matrix, so we approximate the operation by an inner CG solve preconditioned with the classical algebraic multigrid (AMG) method. AMG has proved robust for the irregular sparsity patterns that arise in complex geometries  [56]—such as those produced by topology optimization.
For the Stokes equations, this numerical method—which we refer to as CG-Uzawa—was first proposed in [57], and its superlinear convergence at the operator level for smooth domains was later demonstrated by Axelsson et al. [58]. In the context of porous-media flow, the CG-Uzawa method was implemented and analyzed by Pimanov et al. [49]. To the best of the authors’ knowledge, the present work is the first application of the method to the Stokes–Brinkman equations.
@@ 5 Numerical Experiments
The optimization subproblems are solved with GLOP, the linear programming component of Google OR‑Tools [59], which implements a revised primal–dual simplex algorithm optimized for sparse matrices. Linear systems involving the perturbed velocity Laplacian, $\nu\mathbf{L}_{U}+\mathbf{A}(\Gamma)$, and the heat convection–diffusion operator, $k\mathbf{L}_{\Theta}+\mathbf{C}\bigl{(}U(\Gamma)\bigr{)}$, together with its adjoint, are solved using the algebraic multigrid library SAMG  [60] with the default settings.
@@ 5.1 Example 1: Optimization of a fluid manifold with multiple outlets
In this subsection we revisit the pure-fluid manifold optimization problem with multiple outlets, as studied by Liu et al.  [61]. The objectives are twofold:
@@
For cases (i)–(iii) we employ the CG–Uzawa algorithm described in Section 4.1 with tolerance $\varepsilon_{S}=10^{-3}$. Case (iv) uses the CG–SIMPLE algorithm from  [27]. Convergence histories are plotted in Figure 6, and Table 1 reports the CPU time, RAM usage, and final objective values for all four scenarios. Since the final topologies are visually indistinguishable in all cases, only one representative design is shown (Fig. 5(c)).
@@ 5.2 Example 2: Optimization of a single‑fluid heat exchanger under a geometric constraint
In this section, we optimize a single‑fluid heat exchanger. The setting is similar to that studied by Kambampati et al. [62], except that we model Stokes flow rather than Darcy flow and assume a larger fluid volume fraction. Our aim is to show how the proposed algorithm accommodates additional geometric constraints when only a subset of the domain $\Omega$ is designable. The same treatment of non‑design regions is later used to impose a minimum‑thickness constraint.
@@ 5.3 Example 3: Optimization of a Two‑Fluid Heat Exchanger
In this example, we revisit the two‑fluid heat‑exchanger problem from Feppon et al. [20]. The domain consists of two separate channels, $\Omega_{f}^{\mathrm{hot}}$ and $\Omega_{f}^{\mathrm{cold}}$, that carry different fluids. The solid phase conducts heat between the channels while keeping them apart, ensuring that the two fluids do not interpenetrate.
@@
Figure 10(a) shows the boundary conditions [20]. The hot inlet, $\partial\Omega_{\mathrm{in}}^{\mathrm{hot}}$ (red), is a circle of radius $r=0.1$ on the plane $y=0$, and the hot outlet, $\partial\Omega_{\mathrm{out}}^{\mathrm{hot}}$ (green), lies on $y=1$. The cold inlet, $\partial\Omega_{\mathrm{in}}^{\mathrm{cold}}$ (blue), is located on $x=0$, and the cold outlet, $\partial\Omega_{\mathrm{out}}^{\mathrm{cold}}$ (orange), is on $x=1$.
@@
Because the velocity vanishes on the fluid–solid interface and both velocity and temperature are prescribed at the inlets, maximizing $\Phi_{\theta}(\gamma)$ is equivalent to maximizing the objective used by Feppon et al. [20]:
@@
Compared with the original formulation in [20], we impose the Dirichlet velocity conditions at both the inlets and outlets instead of prescribing outlet pressure. Consequently, the functional $\Phi_{\theta}(\gamma)$ is linear in the state variable $\theta$, which simplifies the sensitivity analysis.
@@
Finally, we model Stokes flow rather than the full Navier–Stokes equations, which greatly simplifies computation of sensitivities. For the Reynolds numbers considered in [20], this simplification introduces a difference in the objectives of less than $1\%$.
@@
In what follows, two scenarios are considered. Section 5.3.1 imposes a non‑mixing constraint, i.e., the minimum‑thickness constraint with a one‑voxel separation. The thermal conductivity is set to $k=5\times 10^{-4}$, which yields a Peclet number $Pe=2000$, matching the value used by Feppon et al. [20]. The domain is discretized with $n=300$ grid points per direction, corresponding to $27\times 10^{6}$ design variables. The goal of this subsection is to examine how the final design, objective value, and convergence depend on the weight $\omega$ in the objective  (63).
@@ 5.3.2 Large-scale validation with COMSOL Multiphysics
6 Conclusions
We have introduced a sparse narrow‑band topology optimization framework for conjugate thermal–fluid problems that (i) restricts design updates to the fluid–solid interface, (ii) eliminates fictitious solid voxels from the flow forward and adjoint analysis, and (iii) utilizes an efficient finite-difference flow solver optimized for voxel-based geometries. In doing so, we retain the flexibility of density methods while approaching the computational efficiency of level‑set techniques.
Three test cases underline the practicality and robustness of the approach. For a $180^{3}$ manifold (Example 1), excluding fictitious solids cuts the wall time from 5.6 h to 22 min and the memory from 26 GB to 4 GB, a $15\times$ and $6.5\times$ reduction, respectively, without altering the final objective. The single‑fluid heat exchanger study (Example 2) confirms that the algorithm converges smoothly even when the objective couples pressure drop and probe point temperature. On a $200^{3}$ grid, each outer iteration costs 32 s for a fluid flow solve and 17 s for a heat transfer solve, and requires less than 17 GB of RAM. Finally, a two‑fluid exchanger with up-to $51\times 10^{6}$ design variables (Example 3) is optimized with a minimal thickness constraint. A complete forward–adjoint cycle for the final design takes $9$ min for the Stokes‑Brinkman system and $2.5$ min for the convection–diffusion system while using less than 130 GB of memory on 12 CPU cores. In addition, the performance predictions deviate by less than 4.3% from an independently meshed COMSOL model, confirming the solver’s accuracy.
These results demonstrate that large‑scale, multiphysics topology optimization—previously confined to distributed memory computers—can be executed on high‑end workstations with turnaround times compatible with industrial design loops.

### v2

@@ 4.1 Numerical method and preconditioning
Because the matrix $\mathbf{S}$ is symmetric positive definite, we solve (30) with the conjugate-gradient (CG) method, which can be viewed as a Krylov acceleration of the classical Uzawa algorithm. Each application of $\mathbf{S}$ requires the action of the inverse of the perturbed velocity Laplacian $\nu\overline{\mathbf{L}}_{U}+\overline{\mathbf{A}}(\Gamma)$. Forming this inverse explicitly would produce a dense matrix, so we approximate the operation by an inner CG solve preconditioned with the classical algebraic multigrid (AMG) method. AMG has proved robust for the irregular sparsity patterns that arise in complex geometries  [ruge1987algebraic]—such as those produced by topology optimization.
For the Stokes equations, this numerical method—which we refer to as CG-Uzawa—was first proposed in [verfurth1984combined], and its superlinear convergence at the operator level for smooth domains was later demonstrated by Axelsson et al. [axelsson2021krylov]. In the context of porous-media flow, the CG-Uzawa method was implemented and analyzed by Pimanov et al. [pimanov2024efficient]. To the best of the authors’ knowledge, the present work is the first application of the method to the Stokes–Brinkman equations.
@@ 5 Numerical Experiments
The optimization subproblems are solved with GLOP, the linear programming component of Google OR‑Tools [ortools_routing], which implements a revised primal–dual simplex algorithm optimized for sparse matrices. Linear systems involving the perturbed velocity Laplacian, $\nu\mathbf{L}_{U}+\mathbf{A}(\Gamma)$, and the heat convection–diffusion operator, $k\mathbf{L}_{\Theta}+\mathbf{C}\bigl(U(\Gamma)\bigr)$, together with its adjoint, are solved using the algebraic multigrid library SAMG  [stuben2017algebraic] with the default settings.
@@ 5.1 Example 1: Optimization of a fluid manifold with multiple outlets
In this subsection we revisit the pure-fluid manifold optimization problem with multiple outlets, as studied by Liu et al.  [liu2022marker]. The objectives are twofold:
@@
For cases (i)–(iii) we employ the CG–Uzawa algorithm described in Section 4.1 with tolerance $\varepsilon_{S}=10^{-3}$. Case (iv) uses the CG–SIMPLE algorithm from  [pimanov2022workflow]. Convergence histories are plotted in Figure 6, and Table 1 reports the CPU time, RAM usage, and final objective values for all four scenarios. Since the final topologies are visually indistinguishable in all cases, only one representative design is shown (Fig. 5(c)).
@@ 5.2 Example 2: Optimization of a single‑fluid heat exchanger under a geometric constraint
In this section, we optimize a single‑fluid heat exchanger. The setting is similar to that studied by Kambampati et al. [kambampati2020level], except that we model Stokes flow rather than Darcy flow and assume a larger fluid volume fraction. Our aim is to show how the proposed algorithm accommodates additional geometric constraints when only a subset of the domain $\Omega$ is designable. The same treatment of non‑design regions is later used to impose a minimum‑thickness constraint.
@@ 5.3 Example 3: Optimization of a Two‑Fluid Heat Exchanger
In this example, we revisit the two‑fluid heat‑exchanger problem from Feppon et al. [feppon2021body]. The domain consists of two separate channels, $\Omega_{f}^{\mathrm{hot}}$ and $\Omega_{f}^{\mathrm{cold}}$, that carry different fluids. The solid phase conducts heat between the channels while keeping them apart, ensuring that the two fluids do not interpenetrate.
@@
Figure 10(a) shows the boundary conditions [feppon2021body]. The hot inlet, $\partial\Omega_{\mathrm{in}}^{\mathrm{hot}}$ (red), is a circle of radius $r=0.1$ on the plane $y=0$, and the hot outlet, $\partial\Omega_{\mathrm{out}}^{\mathrm{hot}}$ (green), lies on $y=1$. The cold inlet, $\partial\Omega_{\mathrm{in}}^{\mathrm{cold}}$ (blue), is located on $x=0$, and the cold outlet, $\partial\Omega_{\mathrm{out}}^{\mathrm{cold}}$ (orange), is on $x=1$.
@@
Because the velocity vanishes on the fluid–solid interface and both velocity and temperature are prescribed at the inlets, maximizing $\Phi_{\theta}(\gamma)$ is equivalent to maximizing the objective used by Feppon et al. [feppon2021body]:
@@
Compared with the original formulation in [feppon2021body], we impose the Dirichlet velocity conditions at both the inlets and outlets instead of prescribing outlet pressure. Consequently, the functional $\Phi_{\theta}(\gamma)$ is linear in the state variable $\theta$, which simplifies the sensitivity analysis.
@@
Finally, we model Stokes flow rather than the full Navier–Stokes equations, which greatly simplifies computation of sensitivities. For the Reynolds numbers considered in [feppon2021body], this simplification introduces a difference in the objectives of less than $1\%$.
@@
In what follows, two scenarios are considered. Section 5.3.1 imposes a non‑mixing constraint, i.e., the minimum‑thickness constraint with a one‑voxel separation. The thermal conductivity is set to $k=5\times 10^{-4}$, which yields a Peclet number $Pe=2000$, matching the value used by Feppon et al. [feppon2021body]. The domain is discretized with $n=300$ grid points per direction, corresponding to $27\times 10^{6}$ design variables. Within this scenario we examine (i) the dependence on the weight $\omega$ in the objective (63) (Section 5.3.1), and (ii) the effect of the initial total fluid volume while fixing the weight $\omega$ (Section 5.3.2).
@@ 5.3.1 Effect of the weight $\omega$
Although the choice $V_{0}^{\mathrm{cold}}=V_{0}^{\mathrm{hot}}=0.15$ (used in part to enable direct comparison with Feppon et al.) implies that a large fraction of the computational bounding box is solid, most of this solid remains outside of the area where conduction and convection are significant. The amount of solid between the phases—the part that matters for conduction—is governed by the minimum‑separation constraint. Starting from well‑separated initial channels, the optimizer systematically moves the phases closer to each other (as close as the prescribed gap allows): The intervening wall becomes thinner and is often perforated by conductive bridges that enhance exchange, rather than forming bulky barriers.
@@
5.3.2 Effect of the initial fluid volume
Since the optimization algorithm preserves the initial volume of each phase, we additionally investigate how the initial total fluid volume influences the outcome. Using the same boundary conditions as in Fig. 10(a), the same Dirichlet velocity profiles (58), and the same non‑mixing constraint as in Section 5.3.1, we fix $\omega=0.4$ and vary the total fluid volume $V_{0}^{\mathrm{cold}}+V_{0}^{\mathrm{hot}}\in\{0.2,\,0.3,\,0.5\}$. Different volumes are realized by adjusting the initial channel thicknesses while keeping inlet/outlet radii fixed. Optimized designs for the extreme cases $0.2$ and $0.5$ are shown in Fig. 13; the  0.3\,0.3 case corresponds to the design shown in Fig. 11(d).
(a) $V_{0}^{\mathrm{cold}}+V_{0}^{\mathrm{hot}}=0.2$
(b) $V_{0}^{\mathrm{cold}}+V_{0}^{\mathrm{hot}}=0.5$
Figure 13: Optimized designs at $\omega=0.4$ for different total initial fluid volumes. The non‑mixing constraint and boundary conditions are identical to those in Section 5.3.1.
Table 6 summarizes the objective values. Since the Dirichlet inlet/outlet velocities are identical in both cases, the net exchanged heat $\Phi_{\theta}^{\text{cold}}-\Phi_{\theta}^{\text{hot}}$ varies only modestly with the total fluid volume. In contrast, the combined pressure drop $\Phi_{(u,p)}^{\text{cold}}+\Phi_{(u,p)}^{\text{hot}}$ increases sharply as the available fluid volume decreases, reflecting the higher viscous losses in thinner channels.
Table 6: Final objective values for $\omega=0.4$ for different total initial fluid volumes.
$V_{0}^{\mathrm{cold}}+V_{0}^{\mathrm{hot}}$ $\Phi_{(u,p)}^{\text{cold}}$ $\Phi_{(u,p)}^{\text{hot}}$ $\Phi_{(u,p)}^{\text{cold}}+\Phi_{(u,p)}^{\text{hot}}$ $\Phi_{\theta}^{\text{cold}}$ $\Phi_{\theta}^{\text{hot}}$ $\Phi_{\theta}^{\text{cold}}-\Phi_{\theta}^{\text{hot}}$
0.2 35.4 65.2 100.6 1.13 0.43 0.69
0.3 13.3 19.7 33.0 1.12 0.45 0.67
0.5 4.1 9.3 13.4 1.07 0.50 0.57
Overall, reducing the initial fluid volume produces thinner and more tortuous channels that preserve heat‑exchange performance but substantially amplify the required pressure drop. Conversely, increasing the fluid volume reduces the hydrodynamic cost with only a modest decrease in net heat removed by the cold fluid.
5.3.3 Dependence on the initial design
For completeness, we also perform optimizations starting from different initial guesses. Following [feppon2021body], our baseline initialization consists of two nested arrays of pipes (see Fig. 10(b)): a $5\times 5$ array for the cold phase and a $4\times 4$ array for the hot phase. Here we consider two additional initial designs: (i) a $3\times 3$ array for the cold phase and a $2\times 2$ array for the hot phase, and (ii) a $7\times 7$ array for the cold phase and a $6\times 6$ array for the hot phase. The corresponding initial total fluid volumes are $V_{0}^{\mathrm{cold}}+V_{0}^{\mathrm{hot}}=0.25$ and $V_{0}^{\mathrm{cold}}+V_{0}^{\mathrm{hot}}=0.40$, respectively. We perform the optimizations with weights $\omega=0.1$ and $\omega=0.2$, respectively. The initial and final designs are shown in Fig. 14.
From these cases we observe a pronounced dependence on the initial design. This is a result of the stiff optimization problem primarily arising from the non-mixing constraint. Accordingly, a reasonable choice of the starting topology can influence both the convergence behavior and the quality of the final design when optimizing two-fluid heat exchangers with non-mixing constraint. In practice, additional manufacturing and space constraints would indicate a good starting topology.
(a) Initial: $3\times 3$ (cold), $2\times 2$ (hot)
(b) Initial: $7\times 7$ (cold), $6\times 6$ (hot)
(c) Final (from $3\times 3$ / $2\times 2$)
(d) Final (from $7\times 7$ / $6\times 6$)
Figure 14: Initial (top) and optimized (bottom) designs for two different initial topologies. The non‑mixing constraint between cold and hot phases is enforced.
@@ 5.3.4 Large-scale validation with COMSOL Multiphysics
6 Conclusions and future work
We have introduced a sparse, narrow‑band topology optimization methodology for conjugate thermal–fluid problems. We have demonstrated that fictitious solid voxels can be eliminated from the forward and adjoint flow analyses. In doing so, we retain the flexibility of density methods while approaching the computational efficiency of the advanced level‑set techniques. Through multiple large‑scale examples, we have shown that the optimization converges to binary designs within the classical Brinkman penalization framework, without introducing filters or projections.
To demonstrate the method’s efficiency, we optimized a two‑fluid exchanger with up to $51\times 10^{6}$ design variables (Example 3) subject to a minimum‑thickness constraint. A complete forward–adjoint cycle for the final design took $9$ min for the Stokes–Brinkman system and $2.5$ min for the convection–diffusion system, while using less than $130$ GB of memory on $12$ CPU cores. In addition, the performance predictions deviate by less than $4.3\%$ from an independently meshed COMSOL model, confirming the solver’s accuracy. These results demonstrate that large‑scale, multiphysics topology optimization—previously confined to distributed‑memory computers—can be executed on high‑end workstations with turnaround times compatible with industrial design loops. As future work, we plan to extend the present Stokes–Brinkman formulation to the Navier– Stokes flow.